Please find the area of a rectangle whose length is 21 cm and whose perimeter is 90 cm.
Answer by dudeantariksh(10)

hiee there..
firstly..
we know that in a rectangle , the opposite sides are equal......................
we also know that area of a rectangle is (length) x (breadth)
so if we find the breadth then we can easily compute the area.
................................................................................
perimeter of any figure is the sum of the lenghts of all sides..
therefore,
perimeter of a rectangle = lenghth + length +breadth + breadth
= 2(length + breadth)
= 2( 21 + breadth)
................................................................................
but perimeter of the rectangle is 90 cm
therefore,
90 = 2(21+breadth)
therefore, if we divide by 2 throughout, we get,
90/2 = 2/2( 21+ breadth)
therefore,
45 = 21 + breadth
therefore, breadth = 45-21 (subtracting 21 from both sides)
breadth = 24 cm.
...................................................................
area of the rectangle = length x breadth
= 21 x 24
= 504 sq. cm.
ans.. the area of the rectangle is 504 sq.cm.